# Rational solutions of Painlevé V from Hankel determinants and the asymptotics of their pole locations

### v2

@@ 4.1 Outside $EoT$: genus 00 $g$–function and effective potential
$\varphi(z;s)=2\ln\left(\frac{2iz}{s}+i\sqrt{\frac{4z^{2}}{s^{2}}+1}\right)-2\sqrt{1+\frac{s^{2}}{4z^{2}}}.$ (4.18)
We need to describe the domain of analyticity; we start from the domain of analyticity of $\varphi^{\prime}$. A great simplification is achieved by observing that $\varphi$ is really a function only of $\frac{z}{s}$, namely
$\varphi(z;s)=\varphi\left(\frac{z}{s};1\right),$ (4.19)
@@
$\varphi_{0}(z)=2\ln\left({2iz}+i\sqrt{{4z^{2}}+1}\right)-2\sqrt{1+\frac{1}{4z^{2}}},\ \ \ \varphi_{0}^{\prime}(z)=\frac{\sqrt{4z^{2}+1}}{z^{2}}.$ (4.20)
@@
The determination of the root is chosen such that $\varphi^{\prime}_{0}(z)\simeq\frac{2}{z}$ at $z=\infty$, with a branch cut connecting the branchpoints ±$\pm\frac{i}{2}$ to be determined below. The language of vertical trajectories of quadratic differentials of [55] is useful in this discussion: by definition these are the arcs of curves where $\mathrm{Re}\,\varphi_{0}$ is constant, which, in the plane of the variable $\xi(z):=\varphi_{0}(z)=\int^{z}\varphi_{0}^{\prime}(w){\mathrm{d}}w$ are (by definition) vertical segments, whence the terminology. We start by observing that $\mathop{\mathrm{res}}_{z=\infty}\varphi^{\prime}_{0}(z){\mathrm{d}}z=-2i\pi$ and $\mathop{\mathrm{res}}_{z=0}\varphi^{\prime}_{0}(z){\mathrm{d}}z=\pm 2i\pi$ (with the sign depending on whether the branch cut leaves $z=0$ to the left or to the right) and hence, no matter how we choose the branch cut $\Gamma_{m}$ (connecting the branchpoints) we have that
the function $\mathrm{Re}\,\varphi_{0}(z)$ is single–valued, harmonic in ${\mathbb{C}}\setminus\Gamma\cup\{0\}$ and continuous in ${\mathbb{C}}\setminus\{0\}$; (4.21)
for $|z|$ sufficiently large $\mathrm{Re}\,\varphi_{0}(z)=\ln|z|+$ harmonic and bounded. (4.22)
The observation (4.22) implies that the level curves of $\mathrm{Re}\,\varphi_{0}$ are deformed circles for $|z|$ sufficiently large. One can verify that changing the determination of both radicals in (4.20) has the effect of flipping the sign of $\mathrm{Re}\,\varphi_{0}$ and hence that $\mathrm{Re}\,\varphi_{0}$ is a well–defined harmonic function on the Riemann surface of the radical
$w^{2}={1+4z^{2}}.$ (4.23)
@@
Furthermore $\mathrm{Re}\,\varphi_{0}$ is an odd function under the holomorphic involution that maps $(w,z)$ to $(-w,z)$. This means that the level sets $\mathrm{Re}\,\varphi_{0}=0$ are well defined on the $z$–plane; they consist of “vertical trajectories” (in the terminology of [55]) issuing from the points ±$\pm\frac{i}{2}$ and forming the pattern illustrated in Fig. 4.
@@
We choose the branch cut of the radical as the arc of Fig. 4 joining ±$\pm\frac{i}{2}$ in the right half–plane. With this choice we have that
$\varphi_{0}^{\prime}(z)\simeq-\frac{1}{z^{2}}+\mathcal{O}(1),\ \ \ z\to 0,$ (4.24)
and in general $\varphi^{\prime}(z;s)=\frac{1}{s}\varphi_{0}^{\prime}\left(\frac{z}{s}\right)$ satisfies
$\varphi^{\prime}(z;s)=-\frac{s}{z^{2}}+\mathcal{O}(1).$ (4.25)
Verification of the properties of $\varphi$ and range of validity.
@@
It suffices to verify the properties for $s=1$ since changing $s\in{\mathbb{C}}$ just amounts to a complex homothethy $z\mapsto sz$. We choose $\Gamma_{m}$ as the arc joining ±$\pm\frac{i}{2}$ in the right plane, $\Gamma_{c}$ as an arc joining the two points ±$\pm\frac{i}{2}$ in the left plane, and inside the region bounded by the imaginary axis and the contour $-\Gamma_{m}$ (see Fig. 4). Finally we choose $\Gamma_{\infty}$ as the ray $(-i\infty,-\frac{i}{2}]$. We then proceed with the verification of the properties in Cor. 4.1:
  1. 1.
@@
  2. 2.
    On $\Gamma_{c}$ we similarly have $\varphi(z_{+})=\varphi(z_{-})$.
  3. 3.
    On $\Gamma_{\infty}$ we have $\varphi_{0}(z_{+})=\varphi_{0}(z_{-})-\mathop{\mathrm{res}}_{w=\infty}\varphi_{0}^{\prime}(w){\mathrm{d}}w=\varphi_{0}(z_{-})+4i\pi$.
  4. 4.
    Since $\Gamma_{m}$ is defined as the zero level set of $\mathrm{Re}\,\varphi_{0}$, we have $\mathrm{Re}\,\varphi_{0}\equiv 0$ on $\Gamma_{m}$ by definition.
  5. 5.
    In the unbounded doubly connected region outside of the “apricot” in Fig. 4 we have $\mathrm{Re}\,\varphi_{0}=\ln|z|+\mathcal{O}(1)$ near $z\to\infty$; thus inevitably $\mathrm{Re}\,\varphi_{0}>0$ in the whole region (which, we remind, is bounded by the zero level sets of $\mathrm{Re}\,\varphi_{0}$).
  6. 6.
    In the right hemi-apricot, the sign must be also positive because $\mathrm{Re}\,\varphi_{0}(z)=\mathrm{Re}\,\left(\frac{1}{z}+\mathcal{O}(1)\right)$.
  7. 7.
@@
Thus all conditions except possibly the condition no. 2 in Def. 4.1 are verified, namely, we still need to verify that the union $\Gamma_{m}\cup\Gamma_{c}$ is homotopic to a circle $|z|=R$, $R>1$ in the cut plane ${\mathbb{C}}\setminus[0,1]$.
Since the level sets in Fig. 4 are scaled by $s$, this latter condition is fulfilled as long as the point $z=1$ lies inside the re-scaled apricot. This holds clearly for $|s|$ sufficiently large, and it fails precisely when the point $z=1$ lies on either $\Gamma_{m}$ or $\Gamma_{c}$, namely when
$\mathrm{Re}\,\varphi(1;s)=0=\mathrm{Re}\,\varphi_{0}\left(\frac{1}{s}\right).$ (4.26)
Observing (4.26) we conclude that the shape of the locus (4.26) is simply the image of the apricot in Fig. 4 (left pane) under the inversion $z\mapsto\frac{1}{z}$ (see right pane in Fig. 4). In Fig. 5 we illustrate the contour for a generic value of $s$ outside the set $EoT$.
@@
Theorem 4.1
For $s$ outside the region $EoT$ the effective potential is given by (4.18) (and (4.17)). The contour $\gamma$ can be chosen as the contour $\Gamma_{m}\cup\Gamma_{c}$ consisting of the arc $\Gamma_{m}$ from $-\frac{is}{2}$ to $\frac{is}{2}$ lying on the left of the straight segment, passing through the origin, from $\frac{is}{2}$ to $-\frac{is}{2}$ (in this orientation). The arc $\Gamma_{c}$ is an arc in the right “lobe” keeping $z=0$ to its left. See Fig. 5.
@@
The only statement that has not been proven yet is the expression (4.27). The expression for $g(z;s)$ is derived from that of the effective potential (4.18) and the relation (4.10) expressing the effective potential in terms of the $g$–function. Since $g(z)=\ln(z)+\mathcal{O}(z^{-1})$ (note the absence of a constant term in the asymptotic expansion) we can deduce $\ell$ by the expansion at $z=\infty$ of $g(z;s)-\ln(z)$.
From (4.18), (4.10) we have
$g(z;s)-\ln z=\ln\left(\frac{2}{s}\right)+\ln\left(-1-\sqrt{1+\frac{s^{2}}{4z^{2}}}\right)-\frac{\ell}{2}+\mathcal{O}(z^{-1})=\ln\left(-\frac{4}{s}\right)-\frac{\ell}{2}+\mathcal{O}(z^{-1}).$ (4.28)
@@
Figure 4: Left: The level sets of $\mathrm{Re}\,\varphi_{0}(z)=0$. Right: the boundary of the region of validity of the genus–zero assumption (ignore the vertical rays issuing from ±$\pm 2i$). It is the locus of $\mathrm{Re}\,\varphi_{0}(1/s)=0$. The inside of this region we refer to as the “Eye of the Tiger” ($EoT$).
Figure 5: The contours $\Gamma_{m},\Gamma_{c}$ and the regions where $\mathrm{Re}\,\varphi(z;s)<0$ (shaded). Indicated also the boundaries, $\mathscr{L}_{\pm}$, of the lens regions $\Lambda_{\pm}$.
4.2 Inside $EoT$: genus one $g$–function and effective potential
The boundary of $EoT$ is precisely the condition that the point $z=1$ belongs to the two sub-arcs of the zero-level set of $\mathrm{Re}\,\varphi(z;s)$ forming the “rind” of the apricot. As we move $s$ inside $EoT$ we cannot use the same effective potential described in the previous section because the second condition in Def. 4.1 ceases to be verified, namely the contour of integration $\gamma$ cannot be homotopically retracted to $\Gamma_{m}\cup\Gamma_{c}$ within ${\mathbb{C}}\setminus[0,1]$ since either $\Gamma_{c}$ or $\Gamma_{m}$ intersect the segment $[0,1]$.
The idea to resolve the impasse is to treat $z=1$ as a “hard–edge”, using the terminology that has come to pass in the random matrix theory literature [9, 12]. We thus postulate the following form for $\varphi^{\prime}(z;s)$
$\varphi^{\prime}(z;s)=\frac{2}{z^{2}}\sqrt{z^{2}+\frac{s^{2}}{4}+\frac{Az^{2}}{z-1}}=\frac{2\sqrt{z^{2}(z-1+A)+\frac{s^{2}}{4}(z-1)}}{z^{2}\sqrt{z-1}}.$ (4.29)
The parameter $A=A(s)$ is chosen by the condition that all periods of $\varphi^{\prime}(z;s){\mathrm{d}}z$ on the Riemann surface of the radical are purely imaginary (this is called Boutroux condition), which is the necessary condition so that $\mathrm{Re}\,\varphi$ is continuous across the cuts; the Riemann surface of $\varphi^{\prime}(z;s)$ is an elliptic curve branched at $z=1$ and the other three roots of the radical in the numerator:
$\mu^{2}=(z-1)\left(z^{2}(z-1+A)+\frac{s^{2}}{4}(z-1)\right).$ (4.30)
We denote these roots as $b,a_{+},a_{-}$ with $b$ the closest root to $z=1$; an expression in terms of Cardano’s formulæ is possible but not necessary.
@@ 4.2 Inside $EoT$: genus one $g$–function and effective potential
$\mathrm{Re}\,\int_{b}^{1}\varphi^{\prime}(z;s){\mathrm{d}}z=0,\qquad\mathrm{Re}\,\int_{b}^{a_{+}}\varphi^{\prime}(z;s){\mathrm{d}}z=0.$ (4.31)
Under these conditions it then follows that the real part of
$\varphi(z;s)=\int_{a_{-}}^{z}\varphi^{\prime}(w;s){\mathrm{d}}w$ (4.32)
@@
By the same argument already used in the genus-zero case, the zero level sets $\mathrm{Re}\,\varphi(z;s)=0$ are well defined; they consist of the critical vertical trajectories of the quadratic differential $Q=\varphi^{\prime}(z;s)^{2}{\mathrm{d}}z^{2}$ [55]
$Q=4\frac{{z^{2}(z-1+A)+\frac{s^{2}}{4}(z-1)}}{z^{4}(z-1)}.$ (4.33)
@@
The following discussion is best followed by referring to Fig. 6, in particular the smaller inset vignettes. The main arcs $\Gamma_{m}$ are sub-arcs of the zero level set of $\mathrm{Re}\,\varphi$ and we need to discuss their qualitative topology before proceeding.
The critical points are the three (generically) simple zeros and the simple pole $z=1$; there are three vertical trajectories that issue from each simple zero, while from the simple pole there is only one. The union of the trajectories is a connected planar graph and the unbounded region is put in conformal equivalence with the punctured unit disk by the map
$\zeta={\rm e}^{-\frac{\varphi(z;s)}{4i\pi}},$ (4.34)
which maps the exterior region into the disk $|\zeta|<1$, with $z=\infty$ mapped to $\zeta=0$. Some observations are in order:
  1. 1.
    The level sets of $\mathrm{Re}\,\varphi$ depend only on $s^{2}$ and they are conjugated if we conjugate $s$.
  2. 2.
    One of the zeros of $\varphi^{\prime}$ is connected by a vertical trajectory to $z=1$; we denote this zero by $z=b$; the other two zeros are one in the upper and one in the lower half plane. We denote them by $a_{\pm}$, respectively.
While the level sets of $\mathrm{Re}\,\varphi$ depend on $s^{2}$ alone, we must choose the branch cuts $\Gamma_{m}$ differently according to the cases $\mathrm{Re}\,s>0$ or $\mathrm{Re}\,s<0$; the reason is that the sign distribution of $\mathrm{Re}\,\varphi$ differs in the two cases. This is seen by the following reasoning:
  • In the outside region $\mathrm{Re}\,\varphi=2\ln|z|+\mathcal{O}(1)$ and hence $\mathrm{Re}\,\varphi>0$;
  • near the origin we must have $\varphi(z;s)=-\frac{s}{z}+\mathcal{O}(1)$ and hence for $\mathrm{Re}\,s>0$ the right “lobe” is where $\mathrm{Re}\,\varphi<0$; vice versa it is the left one if $\mathrm{Re}\,s<0$.
The branch cut $\Gamma_{m}$ is then singled out by the fact that across it the function $\mathrm{Re}\,\varphi$ is continuous but not differentiable, namely, $\mathrm{Re}\,\varphi$ has the same sign (positive) on both sides.
Collecting these observations, we thus have determined that
  1. 1.
    For $\mathrm{Re}\,s>0$ the branch cut $\Gamma_{m}$ consists of the three arcs of the vertical trajectories connecting $z=b,1$ and $z=b,a_{+}$ and $z=b,a_{-}$.
  2. 2.
    For $\mathrm{Re}\,s<0$ the branch cut $\Gamma_{m}$ consists of the two arcs of the vertical trajectories connecting $z=b,1$ and $z=a_{+},a_{-}$ (passing to the left of the origin).
See Fig. 6.
@@
Figure 6: Illustration of the zero level sets of $\mathrm{Re}\,\varphi(z;s)$ for various values of $s$; the inset vignettes have their centre at the value of $s$ to which they correspond.
@@
The specification of the constants that appear for the boundary values of $\varphi$ in Cor. 4.1 will be given in Sections 5.2.1 and 5.2.2 for $\mathrm{Re}\,(s)>0$ and $\mathrm{Re}\,(s)<0$, respectively.

### v3

@@ 3. $(3)$
$\varphi(z;s)=2\ln\left(\frac{2{\rm i}z}{s}+{\rm i}\sqrt{\frac{4z^{2}}{s^{2}}+1}\right)-\frac{s}{z}\sqrt{\frac{4z^{2}}{s^{2}}+1}.$ (4.4)
We need to describe the domain of analyticity. We start from the domain of analyticity of $\varphi^{\prime}$. A great simplification is achieved by observing that $\varphi$ is really a function only of $\frac{z}{s}$, namely
$\varphi(z;s)=\varphi\left(\frac{z}{s};1\right),$
@@
$\varphi_{0}(z)=2\ln\bigl({2{\rm i}z}+{\rm i}\sqrt{{4z^{2}}+1}\bigr)-\frac{\sqrt{4z^{2}+1}}{z},\qquad\varphi_{0}^{\prime}(z)=\frac{\sqrt{4z^{2}+1}}{z^{2}}.$ (4.5)
The determination of the root is chosen such that $\varphi^{\prime}_{0}(z)\simeq\frac{2}{z}$ at $z=\infty$, with a branch cut connecting the branch-points $\pm\frac{\rm i}{2}$ to be determined below.
@@
The language of vertical trajectories of quadratic differentials of [56] is useful in this discussion: by definition these are the arcs of curves where $\operatorname{Re}\varphi_{0}$ is constant, which, in the plane of the variable $\xi(z):=\varphi_{0}(z)=\int^{z}\varphi_{0}^{\prime}(w){\mathrm{d}}w$ are (by definition) vertical segments, whence the terminology. We start by observing that $\mathop{\mathrm{res}}_{z=\infty}\varphi^{\prime}_{0}(z){\mathrm{d}}z=-2{\rm i}\pi$ and $\mathop{\mathrm{res}}_{z=0}\varphi^{\prime}_{0}(z){\mathrm{d}}z=\pm 2{\rm i}\pi$ (with the sign depending on whether the branch cut leaves $z=0$ to the left or to the right) and hence, no matter how we choose the branch cut $\Gamma_{m}$ (connecting the branch-points) we have that the function $\operatorname{Re}\varphi_{0}(z)$ is single-valued, harmonic in ${\mathbb{C}}\setminus\Gamma\cup\{0\}$ and continuous in ${\mathbb{C}}\setminus\{0\}$,
for $|z|$ sufficiently large $\operatorname{Re}\varphi_{0}(z)=\ln|z|+$ harmonic and bounded. (4.6)
@@
The observation (4.6) implies that the level curves of $\operatorname{Re}\varphi_{0}$ are deformed circles for $|z|$ sufficiently large. One can verify that changing the determination of both radicals in (4.5) has the effect of flipping the sign of $\operatorname{Re}\varphi_{0}$ and hence that $\operatorname{Re}\varphi_{0}$ is a well-defined harmonic function on the Riemann surface of the radical $w^{2}={1+4z^{2}}$. Furthermore, $\operatorname{Re}\varphi_{0}$ is an odd function under the holomorphic involution that maps $(w,z)$ to (w,z)(-w,z). This means that the level sets $\operatorname{Re}\varphi_{0}=0$ are well defined on the $z$-plane. They consist of “vertical trajectories” (in the terminology of [56]) issuing from the points $\pm\frac{\rm i}{2}$ and forming the pattern illustrated in Figure 4.
@@
We choose the branch cut of the radical as the arc of Figure 4 joining $\pm\frac{\rm i}{2}$ in the right half-plane. With this choice, we have that $\varphi_{0}^{\prime}(z)\simeq-\frac{1}{z^{2}}+\mathcal{O}(1)$, $z\to 0$, and in general $\varphi^{\prime}(z;s)=\frac{1}{s}\varphi_{0}^{\prime}\left(\frac{z}{s}\right)$ satisfies $\varphi^{\prime}(z;s)=-\frac{s}{z^{2}}+\mathcal{O}(1)$.
@@
Verification of the properties of $\boldsymbol{\varphi}$ and range of validity. It suffices to verify the properties for $s=1$ since changing $s\in{\mathbb{C}}$ just amounts to a complex homothety $z\mapsto sz$. We choose $\Gamma_{m}$ as the arc joining $\pm\frac{\rm i}{2}$ in the right plane, $\Gamma_{c}$ as an arc joining the two points $\pm\frac{\rm i}{2}$ in the left plane, and inside the region bounded by the imaginary axis and the contour $\Gamma_{m}$ (see Figure 4). Finally, we choose $\Gamma_{\infty}$ as the ray $\bigl(-{\rm i}\infty,-\frac{\rm i}{2}\bigr]$. We then proceed with the verification of the properties in Corollary 4.2:
  1. (1)
@@
  2. (2)
    On $\Gamma_{c}$, we similarly have $\varphi(z_{+})=\varphi(z_{-})$.
  3. (3)
    On $\Gamma_{\infty}$, we have $\varphi_{0}(z_{+})=\varphi_{0}(z_{-})-\mathop{\mathrm{res}}_{w=\infty}\varphi_{0}^{\prime}(w){\mathrm{d}}w=\varphi_{0}(z_{-})+4{\rm i}\pi$.
  4. (4)
    Since $\Gamma_{m}$ is defined as the zero level set of $\operatorname{Re}\varphi_{0}$, we have $\operatorname{Re}\varphi_{0}\equiv 0$ on $\Gamma_{m}$ by definition.
  5. (5)
    In the unbounded doubly connected region outside of the “apricot” in Figure 4, we have $\operatorname{Re}\varphi_{0}=\ln|z|+\mathcal{O}(1)$ near $z\to\infty$. Thus inevitably $\operatorname{Re}\varphi_{0}>0$ in the whole region (which, we remind, is bounded by the zero level sets of $\operatorname{Re}\varphi_{0}$).
  6. (5)
    In the right hemi-apricot, the sign must be also positive because $\operatorname{Re}\varphi_{0}(z)=\operatorname{Re}\left(\frac{1}{z}+\mathcal{O}(1)\right)$.
  7. (6)
@@
Thus all conditions except possibly the condition (2) in Definition 4.1 are verified, namely, we still need to verify that the union $\Gamma_{m}\cup\Gamma_{c}$ is homotopic to a circle $|z|=R$, $R>1$ in the cut plane ${\mathbb{C}}\setminus[0,1]$.
Since the level sets in Figure 4 are scaled by $s$, this latter condition is fulfilled as long as the point $z=1$ lies inside the re-scaled apricot. This holds clearly for $|s|$ sufficiently large, and it fails precisely when the point $z=1$ lies on either $\Gamma_{m}$ or $\Gamma_{c}$, namely when
$\operatorname{Re}\varphi(1;s)=0=\operatorname{Re}\varphi_{0}\left(\frac{1}{s}\right).$ (4.7)
Observing (4.7), we conclude that the shape of the locus (4.7) is simply the image of the apricot in Figure 4 (left pane) under the inversion $z\mapsto\frac{1}{z}$ (see right pane in Figure 4). In Figure 5, we illustrate the contour for a generic value of $s$ outside the set $\rm EoT$}. Note that, as a consequence of the branch-cuts in the $z$-plane, the branch-cuts of $\varphi_{0}\left(\frac{1}{s}\right)$ in the $s$-plane now extend towards infinity (vertically) from the points $s=\pm 2{\rm i}$. However, only the imaginary part is discontinuous.
@@
Theorem 4.3.
For $s$ outside the region $\rm EoT$}, the effective potential is given by (4.4) ((and (4.3))). The contour $\gamma$ can be chosen as the contour $\Gamma_{m}\cup\Gamma_{c}$ consisting of the arc $\Gamma_{m}$ from $\frac{{\rm i}s}{2}$ to $\frac{{\rm i}s}{2}$ lying on the left of the straight segment, passing through the origin, from $\frac{{\rm i}s}{2}$ to $\frac{{\rm i}s}{2}$ (in this orientation)). The arc $\Gamma_{c}$ is an arc in the right “lobe” keeping $z=0$ to its left. See Figure 5.
@@
The only statement that has not been proven yet is the expression (4.8). The expression for $g(z;s)$ is derived from that of the effective potential (4.4) and the relation (4.2) expressing the effective potential in terms of the $g$-function. Since $g(z)=\ln(z)+\mathcal{O}\bigl(z^{-1}\bigr)$ (note the absence of a constant term in the asymptotic expansion), we can deduce $\ell$ by the expansion at $z=\infty$ of $g(z;s)-\ln(z)$}.
From (4.4), (4.2), we have
$g(z;s)-\ln z=\ln\left(\frac{2}{s}\right)+\ln\left(-1-\sqrt{1+\frac{s^{2}}{4z^{2}}}\right)-\frac{\ell}{2}+\mathcal{O}\bigl(z^{-1}\bigr)=\ln\left(-\frac{4}{s}\right)-\frac{\ell}{2}+\mathcal{O}\bigl(z^{-1}\bigr).$
@@
Figure 4: Left: the level sets of $\operatorname{Re}\varphi_{0}(z)=0$. Right: the boundary of the region of validity of the genus-zero assumption (ignore the vertical rays issuing from $\pm 2{\rm i}$). It is the locus of $\operatorname{Re}\varphi_{0}(1/s)=0$. The inside of this region we refer to as the “Eye of the Tiger” ($\rm EoT$}).
Figure 5: The contours $\Gamma_{m}$, $\Gamma_{c}$ and the regions where $\operatorname{Re}\varphi(z;s)<0$ (shaded). Indicated also the boundaries, $\mathscr{L}_{\pm}$, of the lens regions $\Lambda_{\pm}$.
4.2 Inside $\boldsymbol{{\rm EoT}}$: Genus one $\boldsymbol{g}$-function and effective potential
The boundary of $\rm EoT$} is precisely the condition that the point $z=1$ belongs to the two sub-arcs of the zero-level set of $\operatorname{Re}\varphi(z;s)$ forming the “rind” of the apricot. As we move $s$ inside $\rm EoT$}, we cannot use the same effective potential described in the previous section because the second condition in Definition 4.1 ceases to be verified, namely the contour of integration $\gamma$ cannot be homotopically retracted to $\Gamma_{m}\cup\Gamma_{c}$ within ${\mathbb{C}}\setminus[0,1]$ since either $\Gamma_{c}$ or $\Gamma_{m}$ intersect the segment $[0,1]$.
The idea to resolve the impasse is to treat $z=1$ as a “hard-edge”, using the terminology that has come to pass in the random matrix theory literature [9, 12]. We thus postulate the following form for $\varphi^{\prime}(z;s)$:
$\varphi^{\prime}(z;s)=\frac{2}{z^{2}}\sqrt{z^{2}+\frac{s^{2}}{4}+\frac{Az^{2}}{z-1}}=\frac{2\sqrt{z^{2}(z-1+A)+\frac{s^{2}}{4}(z-1)}}{z^{2}\sqrt{z-1}}.$
The parameter $A=A(s)$ is chosen by the condition that all periods of $\varphi^{\prime}(z;s){\mathrm{d}}z$ on the Riemann surface of the radical are purely imaginary (this is called Boutroux condition), which is the necessary condition so that $\operatorname{Re}\varphi$ is continuous across the cuts. The Riemann surface of $\varphi^{\prime}(z;s)$ is an elliptic curve branched at $z=1$ and the other three roots of the radical in the numerator:
$\mu^{2}=(z-1)\left(z^{2}(z-1+A)+\frac{s^{2}}{4}(z-1)\right).$
We denote these roots as $b$, $_{+}, aa_{-}$ with $b$ the closest root to $z=1$. An expression in terms of Cardano’s formulæ is possible but not necessary.
@@ 4.2 Inside $\boldsymbol{{\rm EoT}}$: Genus one $\boldsymbol{g}$-function and effective potential
$\operatorname{Re}\int_{b}^{1}\varphi^{\prime}(z;s){\mathrm{d}}z=0,\qquad\operatorname{Re}\int_{b}^{a_{+}}\varphi^{\prime}(z;s){\mathrm{d}}z=0.$ (4.9)
Under these conditions, it then follows that the real part of
$\varphi(z;s)=\int_{a_{-}}^{z}\varphi^{\prime}(w;s){\mathrm{d}}w$
@@
Determination of $\boldsymbol{\Gamma_{m}}$ and $\boldsymbol{\Gamma_{c}}$. By the same argument already used in the genus-zero case, the zero level sets $\operatorname{Re}\varphi(z;s)=0$ are well defined. They consist of the critical vertical trajectories of the quadratic differential $Q=\varphi^{\prime}(z;s)^{2}{\mathrm{d}}z^{2}$ [56]
$Q=4\frac{{z^{2}(z-1+A)+\frac{s^{2}}{4}(z-1)}}{z^{4}(z-1)}.$
@@
The following discussion is best followed by referring to Figure 6, in particular the smaller inset vignettes. The main arcs $\Gamma_{m}$ are sub-arcs of the zero level set of $\operatorname{Re}\varphi$ and we need to discuss their qualitative topology before proceeding.
@@
Figure 6: Illustration of the zero level sets of $\operatorname{Re}\varphi(z;s)$ for various values of $s$; the inset vignettes have their centre at the value of $s$ to which they correspond.
The critical points are the three (generically) simple zeros and the simple pole $z=1$. There are three vertical trajectories that issue from each simple zero, while from the simple pole there is only one. The union of the trajectories is a connected planar graph and the unbounded region is put in conformal equivalence with the punctured unit disk by the map $\zeta={\rm e}^{-\frac{\varphi(z;s)}{4{\rm i}\pi}}$, which maps the exterior region into the disk $|\zeta|<1$, with $z=\infty$ mapped to $\zeta=0$. Some observations are in order:
  1. (1)
    The level sets of $\operatorname{Re}\varphi$ depend only on $s^{2}$ and they are conjugated if we conjugate $s$.
  2. (2)
    One of the zeros of $\varphi^{\prime}$ is connected by a vertical trajectory to $z=1$. We denote this zero by $z=b$. The other two zeros are one in the upper and one in the lower half plane. We denote them by $a_{\pm}$, respectively.
While the level sets of $\operatorname{Re}\varphi$ depend on $s^{2}$ alone, we must choose the branch cuts $\Gamma_{m}$ differently according to the cases $\operatorname{Re}s>0$ or $\operatorname{Re}s<0$. The reason is that the sign distribution of $\operatorname{Re}\varphi$ differs in the two cases. This is seen by the following reasoning:
  • In the outside region $\operatorname{Re}\varphi=2\ln|z|+\mathcal{O}(1)$ and hence $\operatorname{Re}\varphi>0$.
  • Near the origin we must have $\varphi(z;s)=-\frac{s}{z}+\mathcal{O}(1)$ and hence for $\operatorname{Re}s>0$ the right “lobe” is where $\operatorname{Re}\varphi<0$. Vice versa it is the left one if $\operatorname{Re}s<0$.
The branch cut $\Gamma_{m}$ is then singled out by the fact that across it the function $\operatorname{Re}\varphi$ is continuous but not differentiable, namely, $\operatorname{Re}\varphi$ has the same sign (positive) on both sides.
Collecting these observations, we thus have determined that (see Figure 6)
  1. (1)
    For $\operatorname{Re}s>0$, the branch cut $\Gamma_{m}$ consists of the three arcs of the vertical trajectories connecting $z=b,1$ and $z=b,a_{+}$ and $z=b,a_{-}$.
  2. (2)
    For $\operatorname{Re}s<0$, the branch cut $\Gamma_{m}$ consists of the two arcs of the vertical trajectories connecting $z=b,1$ and $z=a_{+},a_{-}$ (passing to the left of the origin).
@@
The specification of the constants that appear for the boundary values of $\varphi$ in Corollary 4.2 will be given in Sections 5.2.1 and 5.2.2 for $\operatorname{Re}(s)>0$ and $\operatorname{Re}(s)<0$, respectively.